OpenStudy (anonymous):

to get a B in math Kai must score an 80 on five tests. Scores on the first 5 tests were 84,77,84 and 79. What is the lowest score she can get on the last test and still get a B?

OpenStudy (anonymous):

do you have any clue on how to solve it?

OpenStudy (anonymous):

Hint: Set up an equation and let x=the lowest score she gets on her last test.

OpenStudy (anonymous):

This is an algebraic expression that you have to come up with that includes averages

OpenStudy (anonymous):

another hint: you have 5 tests, and 4 are known. Somehow you have to relate those to an averaged test score, which includes all 5 tests.

OpenStudy (anonymous):

i'm still confused >.<'

OpenStudy (anonymous):

So an average of the four tests goes as follows: (84+77+84+79)/4

OpenStudy (anonymous):

Hint: They gave you all the scores for the first 4 tests but the last one is unknown and what did I tell u previously to do? Make it equal to x. Now you do know how to calculate an average? If so, I think you should be able to create an equation ow.

OpenStudy (anonymous):

what can you do to include a fifth test in the equation i gave you?

OpenStudy (anonymous):

still confused i don't understand at all >.<

OpenStudy (anonymous):

It would be (84+77+84+79+X)/5=80 Where x is the lowest test score, then you rearrange the equation and solve for X

OpenStudy (anonymous):

ok @heycoa has given you solution which he really shouldn't have, he should have let u attempt to do it urself first but anyway, do you know how to collect like terms and solve for x? because that is the last step.

OpenStudy (anonymous):

@jayds I do know that I basically did the problem for kaisan, but he/she seems frustrated and an example can be very fruitful sometimes.

OpenStudy (anonymous):

i know to collect like terms but i'm confused about solving for x

OpenStudy (anonymous):

bring the 5 over to the RHS and add all the four scores on the LHS first.

OpenStudy (anonymous):

would i multiply by 5 to move it?

OpenStudy (anonymous):

yes u are right because u applied the opposite operation.

OpenStudy (anonymous):

when i multiply would the 5 on the LHS cancel?

OpenStudy (anonymous):

What you do to one side you must always do to the other. so there will be a 5 multiplied on both sides, do you know what 5/5 is?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yes u are right, that is exactly what we are doing, multiplying both sides by 5, but because 5/5=1 on the LHS it cancels and becomes nothing but on RHS we still have to times 5 since we are left with that.

OpenStudy (anonymous):

\[5 \times \frac{84+77+84+79+x}{5} = 80 \times 5\] \[Sum of numbers+x=400\]

OpenStudy (anonymous):

since the 5x5 cancels on the LHS we get the bottom line. Sum of the numbers + x = 400

OpenStudy (anonymous):

then u can solve for x.

OpenStudy (anonymous):

and tell me what u get.

OpenStudy (anonymous):

i got x=76

OpenStudy (anonymous):

yes u are right.

OpenStudy (anonymous):

i still don't understand why you had 84+77+84+79+x/5

OpenStudy (anonymous):

that's what u do to calculate an average/mean!!

OpenStudy (anonymous):

same concept as the link I sent u.

OpenStudy (anonymous):

oh! ok so since the 5th test was unknown you put an x in place of it?

OpenStudy (anonymous):

say I want to find the mean of 1,2,3 then I have to add them altogether and divide by 3. Mean = (1+2+3) / 2 = 3

OpenStudy (anonymous):

yes, that is exactly right.

OpenStudy (anonymous):

oh ok ^^' thanks

OpenStudy (anonymous):

no problem.